. If N represents the number of atoms of a radionuclide in a sample at any given time, then the change dN in the number during a short time dt is proportional to N and to d Letting λ be the constant of proportionality (called the decay constant), we write: dN=-ANdt a. Why do we need the negative sign in this equation? b. What are the units of λ? c. From this equation, derive the exponential radioactive decay law/N = N, e 8. The equation derived in question 7 can also be used to calculate activity. Typically, N and No are replaced with A and Aa. Using the equation derived in question 7, calculate the activity of a 30 MBg Na- 22 sample after 2.5 years.


solve number 8
Answer #1

Given that

initial activity Ao=30 MBq

time t=2.5*years=2.5*365=912.5 days

half life T1/2=15 days

now we find the decay constant

decay constant \lambda=0.693/15=0.0462/days

now we find the activity

activity A=Aoe-\lambdat

=30*e^-0.0462*912.5

=1.5 MBq